What is the oxidation number of sulfur in (S₂O₈)²⁻


1
Expert's answer
2022-01-07T09:35:05-0500

Solution:

Structure of S2O82-:



There are two oxygen atoms in the structure that are linked to each through a peroxide bond. The oxidation number of an oxygen atom linked through a peroxide bond is −1.

The remaining oxygen atoms linked to sulphur have an oxidation number of −2.

The total molecule has a charge of −2.

Thus,

(−2) × (6 atoms of O) + (−1) × (2 atoms of O in peroxide linkage) + (x) × (2 atoms of S) = −2

(−2 × 6 ) + (−1 × 2 ) + 2x = −2

−14 + 2x = −2

2x = 12

x = 6

so the oxidation number of sulfur in S2O82- is +6


Answer: The oxidation number of sulfur in S2O82- is +6.